OpenStudy (anonymous):

Please help solve this exponential equation. I used A=P(1+r/n)^nt 80,000=20,000(1+0.06/12)^12(23)

OpenStudy (anonymous):

I really want to see the steps..as this is just a sample problem for me to look at

OpenStudy (campbell_st):

well there are no unknowns in this question... its just a case of evaluating. it can be simply typed into a calculator as its written...

OpenStudy (campbell_st):

the only thing you really need to do is calculate the power... which is 12 x 23 = 276 so enter 20000(1 + 0.06/12)^276

OpenStudy (anonymous):

ok one second

OpenStudy (anonymous):

that doesnt seem right though...

OpenStudy (anonymous):

here is the orignal sample problem Using the same interest, plus the number of years calculated from problem 1, how much principal would the couple need to invest fund the entire cost of college ($80,000)? from the function above we know the interest is 6% and the years is 23.

OpenStudy (anonymous):

solving the equation i get like 80,000

OpenStudy (anonymous):

i am working on log...im assuming i take the log of both sides somewhere...

OpenStudy (campbell_st):

ok..so A the future value is 80000 P = inknown r = 6% p.a. compounded monthly t = 23 years... is that the information..?

OpenStudy (anonymous):

P is 20,000. Thats what the couple invested in the first problem.

OpenStudy (anonymous):

so from this im assuming the equation is set up as so..80,000=20, 000(1+0.06/12)^12(36)

OpenStudy (campbell_st):

so the equation is \[80000 = P(1 + 0.06/12 )^{12\times23}\] then by making P the subject \[P = \frac{80000}{(1 + 0.06/100)^{276}}\]

OpenStudy (campbell_st):

I think the question is asking how much to invest P to have 80000 I think Question 1 would have been what is the future value of the when $20000 is invested at...etc

OpenStudy (anonymous):

here is problem 1... A married couple plan to send their only child to college. They calculate that the first year of college will cost $20,000. If the couple invests $5000 at 6% interest, in how many years will they have at least $20,000 saved? this i solved fine and got 23 years...

OpenStudy (campbell_st):

quick question... does it say anything about interest compounded monthly or yearly..?

OpenStudy (anonymous):

nothing at all

OpenStudy (campbell_st):

ok so assume annually... so then the equation is \[20000 = 5000(1 + 0.06)^n\] divide both sides by 5000 \[4 = (1.06)^n \] using logs base 10 \[\log 4 = n \times \log(1.06)\] divide both sides by log(1.06) n = 23.79 years

OpenStudy (campbell_st):

say 24 years...

OpenStudy (anonymous):

ok so while you were doing this....i came up with this for my answer on #2....$20,202.02

OpenStudy (campbell_st):

then to find the principal to cover $80000 in costs its as its shown above...but I'd assume the interest is calculated annually \[80000 = P(1 + 0.06)^{23.8}\] make P the subject \[P = \frac{80000}{(1.06)^{23.8}}\] which means you need to invest $20915.78...

OpenStudy (campbell_st):

so the amount to invest is between 20000 and 21000 but it depends on the compounding period... if its monthly you will need a slightly smaller amount..

OpenStudy (campbell_st):

I have to go... I hope I've helped and not confused... good luck

OpenStudy (anonymous):

Ok i will ask my teacher tomorrow and see what he says....I really hate these log's though! and yes you did help!